40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 19 12:09:41 2011, josephb, Update, CDS, AA filter box pulled from 1X5, optic suspensions currently off 
    Reply  Thu May 19 17:53:03 2011, steve, Update, CDS, AA filter box modified at 1X5 P1070727.JPG
Message ID: 4750     Entry time: Thu May 19 17:53:03 2011     In reply to: 4748
Author: steve 
Type: Update 
Category: CDS 
Subject: AA filter box modified at 1X5 

Quote:

[Steve, Joe]

Steve pulled the top AA filter box from 1X5 which handled some of the suspensions channels.  We turned off all the watchdogs before pulling it out, as well as recorded which cables were connected to which inputs.

The case  is undergoing a structural modification to have the ADC adapter card which previously was loosely connected via cables, securely attached to the case.

Steve still wants to do some cabling in the rack while the box is out, and will return it this afternoon once he has finished that.

 Job is done. Sus damping are back on. Cabling-strain reliefing are  not finished yet at 1X5 and 1X4

Attachment 1: P1070727.JPG  2.856 MB  | Hide | Hide all
P1070727.JPG
ELOG V3.1.3-